# Multi-Cali Anything: Dense Feature Multi-Frame Structure-from-Motion for Large-Scale Camera Array Calibration

### v1

@@ I INTRODUCTION
Recent advances in Structure-from-Motion (SfM) pipelines, such as Pixel-Perfect SfM [1] and VGGSfM [2], have provided efficient tools for joint estimation of camera parameters and 3D scene structure. These methods can perform camera calibration directly from scene data without a separate calibration session. However, SfM pipelines are typically designed to estimate extrinsics and intrinsics simultaneously, and they do not provide an option to utilize known ground-truth extrinsics and refine only intrinsics, leading to suboptimal intrinsics calibration. In large-scale camera array setups where the extrinsics are already well-known, this lack of flexibility results in unnecessary estimation errors and reduced accuracy. Besides, the single frame processing of these pipelines causes inconsistencies in multi-frame datasets, reducing accuracy and robustness.
@@ III-C Dense Feature Reprojection
To address this issue, we are inspired by Pixel-Perfect SfM [1], which mitigates keypoint noise by performing optimization in feature space rather than relying solely on raw reprojection error. Specifically, for the $i$-th frame and $j$-th camera, we use a convolutional neural network (CNN) to compute a dense feature map $\mathbf{F}_{i,j}$ of size $H\times W\times 128$, where $H$ and $W$ are the height and width of the input image, and the feature map has $128$ channels. The CNN is trained to enforce viewpoint consistency: for a given interest point in the scene, the CNN generates identical feature vectors at its projected locations in images, regardless of the observing cameras. This property allows the reprojection error to be measured not only in image space but also in feature space, effectively reducing the influence of noisy keypoints. After extracting the dense feature map $\mathbf{F}_{i,j}$, we compute a reference feature vector $\mathbf{f}_{\mathbf{x}}$ for each 3D point $\mathbf{x}$. The reference feature is defined as the keypoint feature vector closest to the mean of all keypoint feature vectors within the track:
@@
In practice, it is impossible to load all dense feature maps in memory at the same time, and is very inefficient to perform bicubic interpolation $\mathbf{F}_{i,j}(\cdot)$ and compute losses $\rho(\cdot)$ in the $128$-dimensional feature space. To address this issue, inspired by Pixel-Perfect SfM [1], we only keep $16\times 16$ patches around keypoints, and preprocess these feature patches into $3$-dimensional cost maps:
@@ III-E Implementation
In our implementation, we use $\lambda_{0}=1.0,\lambda_{1}=\lambda_{2}=0.01,\lambda_{3}=0.01,\lambda_{4}=\lambda_{5}=0.02$ in the initialization stage. During optimization, $\lambda_{1},\lambda_{2},\lambda_{4},\lambda_{5}$ are multiplied with a scale factor of $2.0$ after each iteration. The whole process terminates when $\lambda_{1}$ exceeds the threshold $\theta=1\times 10^{6}$. The overall optimization is performed using the Ceres Solver [24], which offers robust and efficient performance for large-scale non-linear optimization tasks.
@@ IV-A Metrics
Figure 5: Comparison of DUSt3R-reconstructed models vs. ground-truth model. First is ground-truth model; the next four are DUSt3R reconstructions with extrinsics fixed to ground-truth values: Linear head (our refined intrinsics), DPT head (our refined intrinsics), Linear head (DUSt3R-estimated intrinsics), DPT head (DUSt3R-estimated intrinsics). Models are visualized in the same coordinate system. DUSt3R uses learning-based method to compute point maps and camera parameters, which results in large distortions.
@@ IV-B Quantitative Results
We compared our approach with several state-of-the-art methods, including COLMAP, Pixel-Perfect SfM, DUSt3R, and VGGSfM, on eight frames of the E057 expression of the Multiface dataset. For COLMAP, Pixel-Perfect SfM, and VGGSfM, we append a BA stage with fixed extrinsics after them to evaluate their performances when ground-truth extrinsics are known. For DUSt3R, its implementation already provides the option to fix extrinsics. For our method, we use the sparse output of Pixel-Perfect SfM with SuperPoint and SuperGlue as our input data.
Table I report absolute and relative errors for focal lengths and principal points. Our multi-frame approach achieved the lowest errors for focal lengths, with $\mathrm{focal}_{\mathrm{abs,mean}}$ of $5.405$ and $\mathrm{focal}_{\mathrm{rel,mean}}$ of $0.712$‰. If multi-frame optimization is disabled, our method still surpasses others, with $\mathrm{focal}_{\mathrm{abs,mean}}$ of $6.598$ and $\mathrm{focal}_{\mathrm{rel,mean}}$ of $0.870$‰. For principal points, our multi-frame approach achieved $1.994$ for $\mathrm{pp}_{\mathrm{abs,mean}}$ and $1.335$‰  for $\mathrm{pp}_{\mathrm{rel,mean}}$, outperforming others except for COLMAP. However, the differences are ignorable, and COLMAP has significantly higher focal length errors. Because camera intrinsics are defined by both focal lengths and principal points, the overall intrinsics of COLMAP are still very inaccurate and can lead to inconsistency in downstream reconstruction tasks.
@@ IV-C Qualitative Results
Fig. 2 visualizes reprojection errors for intrinsics estimated by COLMAP, Pixel-Perfect SfM, VGGSfM, and our method. We sample a few points on the ground-truth mesh and project them onto images using both ground-truth and estimated intrinsics. Our method achieves the smallest reprojection error.
Fig. 3 compares multi-view stereo (MVS) reconstructions using COLMAP’s MVS pipeline with different intrinsics as input. We compute point-wise distances to the ground-truth mesh and visualize them in RGB colors, where blue represents positive errors, red indicates negative errors, and green denotes near-zero deviation. Our method exhibits a tighter concentration around green, demonstrating superior intrinsic accuracy compared to other methods.
@@ IV-D Ablation Study
We conducted an ablation study to evaluate the impact of each component, as summarized in Table II. Starting with the baseline using only the reprojection term (Row 1), the absolute focal length and principal point errors are $7.307$ and $2.391$, respectively. Adding the extrinsics regularization term (Row 2) slightly increases the focal length error, highlighting the importance of progressive coefficient adjustment. With this adjustment (Row 3), errors decrease to $6.616$ and $2.241$. Including the dense feature reprojection term (Row 6) further reduces errors to $6.598$ and $2.227$. Finally, adding the intrinsics variance term for multi-frame optimization (Row 7) achieves the best results, with absolute focal length error of $5.405$ and absolute principal point error of $1.994$, showcasing the combined effectiveness of these components.

### v2

@@ I INTRODUCTION
Recent advances in Structure-from-Motion (SfM) methods, such as Pixel-Perfect SfM [1] and VGGSfM [2], have provided efficient tools for joint estimation of camera parameters and scene structure. These methods can perform camera calibration directly from scene data without separate calibration sessions. However, SfM pipelines are typically designed to estimate extrinsics and intrinsics simultaneously, and do not provide an option to utilize known ground-truth extrinsics and refine only intrinsics, leading to suboptimal intrinsics calibration. In large-scale camera array setups where the extrinsics are already well-known, this lack of flexibility results in unnecessary estimation errors and reduced accuracy. Besides, the single frame processing of these pipelines causes inconsistencies in multi-frame datasets, reducing robustness.
@@ III-C Dense Feature Reprojection
To address this issue, we are inspired by Pixel-Perfect SfM [1], which mitigates keypoint noise by performing optimization in feature space rather than relying solely on raw reprojection error. Specifically, for the $i$-th frame and $j$-th camera, we use a convolutional neural network (CNN) to compute a dense feature map $\mathbf{F}_{i,j}$ of size $H\times W\times 128$, where $H$ and $W$ are the height and width of the input image, and the feature map has 128128128 channels. The CNN is trained to enforce viewpoint consistency: for a given interest point in the scene, the CNN generates identical feature vectors at its projected locations in images, regardless of the observing cameras. This property allows the reprojection error to be measured not only in image space but also in feature space, effectively reducing the influence of noisy keypoints. After extracting the dense feature map $\mathbf{F}_{i,j}$, we compute a reference feature vector $\mathbf{f}_{\mathbf{x}}$ for each 3D point $\mathbf{x}$. The reference feature is defined as the keypoint feature vector closest to the mean of all keypoint feature vectors within the track:
@@
In practice, it is impossible to load all dense feature maps in memory at the same time, and is very inefficient to perform bicubic interpolation $\mathbf{F}_{i,j}(\cdot)$ and compute losses $\rho(\cdot)$ in the 128128128-dimensional feature space. To address this issue, inspired by Pixel-Perfect SfM [1], we only keep $\times 16$ 16 patches around keypoints, and preprocess these feature patches into 333-dimensional cost maps:
@@ III-E Implementation
In our implementation, we use $\lambda_{0}=1.0,\lambda_{1}=\lambda_{2}=0.01,\lambda_{3}=0.01,\lambda_{4}=\lambda_{5}=0.02$ in the initialization stage. During optimization, $\lambda_{1},\lambda_{2},\lambda_{4},\lambda_{5}$ are multiplied with a scale factor of 2.02.02.0 after each iteration. The whole process terminates when $\lambda_{1}$ exceeds the threshold $\theta=1\times 10^{6}$. The overall optimization is performed using the Ceres Solver [24], which offers robust and efficient performance for large-scale non-linear optimization tasks.
@@ IV-A Metrics
Figure 5: Comparison of DUSt3R-reconstructed models vs. ground-truth. First is ground-truth; the next four are DUSt3R reconstructions with extrinsics fixed to ground-truth values: Linear head (our refined intrinsics), DPT head (our refined intrinsics), Linear head (DUSt3R-estimated intrinsics), DPT head (DUSt3R-estimated intrinsics). Models are visualized in the same coordinate system. DUSt3R’s learning-based method to compute point maps results in large distortions.
@@ IV-B Quantitative Results
We compared our approach with several state-of-the-art methods, including COLMAP, Pixel-Perfect SfM, DUSt3R, and VGGSfM, on eight frames of the E057 expression of the Multiface dataset. For COLMAP, Pixel-Perfect SfM, and VGGSfM, we append a BA stage with fixed extrinsics to evaluate their performances when ground-truth extrinsics are known. For DUSt3R, it already provides the option to fix extrinsics. For our method, we use the sparse output of Pixel-Perfect SfM with SuperPoint and SuperGlue as our input.
Table I report absolute and relative errors for focal lengths and principal points. Our multi-frame approach achieved the lowest errors for focal lengths, with $\mathrm{focal}_{\mathrm{abs,mean}}$ of 5.4055.4055.405 and $\mathrm{focal}_{\mathrm{rel,mean}}$ of 0.7120.7120.712‰. If multi-frame optimization is disabled, our method still surpasses others, with $\mathrm{focal}_{\mathrm{abs,mean}}$ of 6.5986.5986.598 and $\mathrm{focal}_{\mathrm{rel,mean}}$ of 0.8700.8700.870‰. For principal points, our multi-frame approach achieved 1.9941.9941.994 for $\mathrm{pp}_{\mathrm{abs,mean}}$ and 1.3351.3351.335‰  for $\mathrm{pp}_{\mathrm{rel,mean}}$, outperforming others except for COLMAP. However, the differences are ignorable, and COLMAP has significantly higher focal length errors. Because camera intrinsics are defined by both focal lengths and principal points, the overall intrinsics of COLMAP are still very inaccurate and can lead to inconsistency in downstream reconstruction tasks.
@@ IV-C Qualitative Results
Fig. 2 visualizes reprojection errors for intrinsics estimated by COLMAP, Pixel-Perfect SfM, VGGSfM, and ours. We sample a few points on ground-truth mesh and project them onto images using both ground-truth and estimated intrinsics. Our method achieves the smallest reprojection error.
Fig. 3 compares multi-view stereo (MVS) reconstructions using COLMAP’s MVS pipeline with different intrinsics as input. We compute point-wise distances to the ground-truth mesh and visualize them in RGB colors, where green denotes near-zero deviation. Our method exhibits a tighter concentration around green, demonstrating superior intrinsic accuracy compared to other methods.
@@ IV-D Ablation Study
We conducted an ablation study to evaluate the impact of each component, as summarized in Table II. Starting with the baseline using only the reprojection term (Row 1), the absolute focal length and principal point errors are 7.3077.3077.307 and 2.3912.3912.391, respectively. Adding the extrinsics regularization term (Row 2) slightly increases the focal length error, highlighting the importance of progressive coefficient adjustment. With this adjustment (Row 3), errors decrease to 6.6166.6166.616 and 2.2412.2412.241. Including the dense feature reprojection term (Row 6) further reduces errors to 6.5986.5986.598 and 2.2272.2272.227. Finally, adding the intrinsics variance term for multi-frame optimization (Row 7) achieves the best results, with absolute focal length error of 5.4055.4055.405 and absolute principal point error of 1.9941.9941.994, showcasing the combined effectiveness of these components.